OpenStudy (anonymous):

what is a three step equation in which the answer is 12?

OpenStudy (anonymous):

equation: whatever you do to one side you must do to the other to ensure equality work from what you are given x = 12 step 1) x + 2 = 14 (add 2 to both sides) step 2) 2x + 4 = 28 (multiply both sides by 2) step 3) (2x + 4)/7 = 4 (divide both sides by 7)